# Building Better: Avoiding Pitfalls in Developing Language Resources when Data is Scarce

### v3

@@ 4.3 Credit Attribution
We asked the respondents to share whether they were properly credited for their work by, e.g., getting financial compensation for a long annotation task, getting involved in the writing of a research paper for a resource that they built, etc. As shown in Figure 5, most respondents (>67%) report this not being the case at least once.
Figure 6 shows the distributions of responses pertaining to how the respondents were incentivised to perform an annotation task for which they were eventually not given due credit.
@@
Figure 6: Respondents on incentives when no proper credit (e.g., financial compensation for data annotation) was offered.
For the respondents who reported that they did not receive proper credit for at least one project they were involved in, we report the initial incentives for joining these projects. As shown in Figure 6, they were either:
@@
    a member of a community (see paragraph below), or
@@
    somehow manipulated.
@@
When using standard crowdsourcing platforms such as AMT or Prolific, one can operationalise the annotation so that it performs better for a given task. Despite their shortcomings Fort et al. (2011); Irani (2015), one can attempt to protect workers by using tests and training when annotating hard tasks. However, for mid- to low-resource languages, platforms such as AMT and Prolific often do not have enough speakers registered on the platform. Therefore, researchers opt for personal connections or community efforts instead. There are various advantages to personal outreach and community efforts, such as the fact that people feel more included and trust can be built more easily. On the other hand, there is a high risk of exploitation and emotional manipulation in such a case, e.g., junior researchers told that being part of an online community that helps build resources for a language is prestigious and worth adding to their CV. For instance, 40% of the respondents, who spent 1 day to more than a month annotating data (see Figure 6), report negative experiences. That is, their work was not properly compensated, acknowledged, or recognised. This calls for a need to set guidelines and standards when using community services.
@@ 5 Recommendations
While there has been a considerable amount of work on the ethics of best practices for building NLP artefacts Bender and Friedman (2018); Mohammad (2022, 2023), our findings substantiate the fact that research on mid- to low-resource languages presents additional challenges that ought to be addressed. This is because it often relies on unconventional practices such as community work and is motivated by more than research challenges and scientific interests as shown in Section 4. Based on our findings, we provide the following recommendations.
@@ 5.1 Center the People
Our findings show that there are various issues that have to be addressed early as research in the area lacks established standards and is subject to power differentials. Many mid- to low-resource languages are from what is called “the Global South” with a large number of them being spoken rather than written.
@@
Language is an important part of a population’s identity and technologies dealing with it have a direct impact on people’s lives. Past NLP work highlights how to engage with speakers and communities whose languages are in question Bird and Yibarbuk (2024); Bird (2020, 2024); Cooper et al. (2024). We further reinforce this argument with our findings.
When a researcher reaches out to a group with little background knowledge of their culture or language, one needs to approach these problems from the perspective of the community in question Bird (2022). Hence, the question of who is exactly served needs to be addressed early to avoid any misconception of perceived needs for language technologies.
@@
In addition to the large percentage of our survey respondents who reported not being properly credited for their labor, there were cases of emotional manipulation (e.g., making emotional arguments such as how one’s labor will help the speakers of the language and that is compensation enough). One has to set rules and expectations with clear communication on the purpose of a given research project. For instance, when dealing with online communities for data collection and annotation, extra care needs to be shown and benevolent prejudice such as depicting oneself as a savior of a local community Bird (2022) ought to be avoided. Companies and research labs relying on communities for annotation and data creation need to properly compensate the contributors.
The question of who is annotating what has to also be addressed. The scarcity of annotators can lead to poor choices as very often, native speakers cannot be found online easily which has led to researchers choosing people from associated regions—people who do not necessarily speak the language variant in question. This results in a problematic overgeneralisation that puts different languages under the same umbrella simply because they have one or a small set of attributes in common. This often leads to potentially oversimplistic solutions. For instance, variations of Arabic differ considerably but numerous research projects have treated entire regions, such as North Africa, as a monolith (e.g., to appear to have more data).
@@ 5.2 Be Fair: Give Credit where Credit is Due
Our findings show the unfortunate trend of data workers and NLP practitioners suffering from a lack of recognition, especially those who are part of online communities that focus on low-resource languages. A needed follow-up work would be extensive fieldwork with the various online communities. Hence, our recommendation is a call to action on the setup of fair and comprehensive practices when collaborating with online communities, while taking power differentials into account. That is, existing authorship standards222 https://www.icmje.org/recommendations/browse/roles-and-responsibilities/defining-the-role-of-authors-and-contributors.html and https://www.aclweb.org/adminwiki/index.php/Authorship_Changes_Policy_for_ACL_Conference_Papers need to be followed and discussed prior to the start of a project as to whether a data worker should be listed as an author. This is particularly critical for junior researchers who substantially contribute to resource construction. Moreover, proper financial compensation needs to be provided for annotators who are essential to the construction of large-scale resources. Ideally, a resource paper should provide proof that the annotators were paid and treated fairly if requested by reviewers as recommended by Rogers et al. (2021).
@@ 5.3 Choose the Jargon Carefully and Be Aware of False Generalisations
As previously discussed in 5.1, it is important to embrace social awareness and avoid grouping people from colonial and Western perspectives Bird (2020, 2022); Held et al. (2023). In this area, we could benefit from critical work in other fields. Hence, one can avoid dismissive and outdated terms and classifications, e.g., “the rest of the world". Note also that the The World’s Values Survey classification Haerpfer and Kizilova (2012) presents an orientalist view of the world Said (1977). It has clear flaws such as including Christian-majority countries (e.g., Ethiopia, Rwanda) in a so-called "African-Islamic" category as well as countries that have very little to do with each other (e.g., Kyrgyzstan and Tunisia). Santy et al. (2023) propose to annotate sentences while taking the culture into consideration. Their approach is sound, however, their classification includes cultures that are too far from each other. For instance, considering Bolivia and the Philippines, or South Africa and UAE as one monolith could create an overgeneralisation or misrepresentation of these cultures.
We further recommend:
  • checking the adequacy of the data with respect to the task (e.g., genre, style, and sources),
  • checking when the data was collected,
  • checking if the data includes offensive utterances,
  • controlling for overgeneralisations and bias.
@@ 5.4 Set Fair and Realistic Expectations
As pointed out by Doğruöz and Sitaram (2022), tools for low-resource languages are often perceived as scaled-down versions of high-resource ones. Adding to previous work elaborating on what this may mean to the speakers Bird (2022); Markl et al. (2024), we focus on the impact of setting these expectations for researchers and practitioners working on mid- to low-resource languages. That is, they may be expected to build models similar to those built for high-resource languages, i.e., tackling the same NLP tasks, and performing extremely well. However, this can be unrealistic for various reasons such as the users’ needs Blaschke et al. (2024), the language’s specific features Bird and Yibarbuk (2024), and the lack of funding linked to the “prestige” of the language as reported by our respondents.
@@
Joshi et al. (2020) survey the state of NLP in various languages. In fact, people do not necessarily need the tools that researchers think they need. Simultaneously, we should not be limiting what NLP research on mid to low-resource languages should be about. This is linked to the focus on local communities as this further reinforces the need to communicate with them Bird and Yibarbuk (2024); Lent et al. (2022); Mager et al. (2023); Cooper et al. (2024).
@@
Researchers dealing with what is considered a “solved problem" for high-resource settings does not mean the research problem is solved. For instance, a language may show properties that distinguish it from what is currently available for English or other high-resource languages, e.g., a rich morphology, the presence of tones Adebara and Abdul-Mageed (2022). Therefore, it is different from what is frequently called “a replication”.
@@ 5.5 Check the Source Even if the Language is Low-resource
Due to the limited amount of online data available for mid- to low-resource languages, there is a tendency to use any online sources to build a tool or a dataset for these languages without looking into the ethical implications of this choice or considering the appropriateness of the data. For instance, it is typically easier to use religious texts, lyrics, or movie subtitles for training a model. However, these should be carefully considered before use Hutchinson (2024); Mager et al. (2023). That is, lyrics are not representative of how people communicate in their daily lives Mayer et al. (2008) since they often rhyme, may contain slang, etc. In addition, Mager et al. (2023) reported concerns about misrepresentations due to the use of religious texts without a thorough inspection of potential implications. Further, we often turn into synthetic data generated using machine translation and using LLMs when these show clear limitations, especially in multicultural settings Hershcovich et al. (2022). Research from other disciplines, even tightly related such as linguistics Turner (2023), is often overlooked. Lessons learned from these disciplines include choosing adequate and suitable data sources as the data will stay for a long time.
@@ 6 Conclusion
We present findings and result analysis of survey responses from NLP researchers and practitioners working on various under-served languages. The responses include common limitations and research practices in the field. We further provide recommendations on how to address the reported limitations while remaining fair to data workers.
Our work is the first to document NLP researchers and workers’ experiences.
7 Limitations
@@ 9 Acknowledgement
Thanks to Michael Schlichtkrull for feedback and discussions on the survey.
@@ A.2 Languages
The full list of the languages that our respondents have worked is included in the following. Note that participants could work on more than one language.

### v5

@@ 4.3 Credit Attribution
We asked the respondents to share whether they were properly credited for their work by, e.g., getting financial compensation for a long annotation task, getting involved in the writing of a research paper for a resource that they built, etc. As shown in Figure 5, most respondents (>67%) report this not being the case at least once. Figure 6 shows the distributions of responses pertaining to how the respondents were incentivised to perform an annotation task for which they were eventually not given due credit.
@@
Figure 6: Respondents on incentives when no proper credit (e.g., financial compensation for data annotation) was offered. We show the counts of the incentives and the time it took participants to complete their work for a given project (from <=2 hours to more than a month).
For the respondents who reported that they did not receive proper credit for projects in which they were involved, we report the incentives for joining these projects and the time it took the participants to complete the work. As shown in Figure 6, they were either:
@@
    a member of a community or a participatory framework (see paragraph below),
@@
    somehow manipulated into thinking that there was a professional benefit in joining without proper compensation as we explain below.
@@
When using standard crowdsourcing platforms such as AMT or Prolific, one can operationalise the annotation for a given task. Despite their shortcomings Fort et al. (2011); Irani (2015), one can attempt to protect workers by using tests and training when annotating hard tasks. However, for mid- to low-resource languages, platforms such as AMT and Prolific often do not have enough speakers registered on the platform. Therefore, researchers opt for personal connections or participatory efforts instead. On the one hand, personal outreach and community efforts make people feel more included and trust can be built more easily. On the other hand, there is a high risk of exploitation and emotional manipulation in such a case. For instance, some junior researchers reported cases of exploitation due to them being told that joining an online community that builds language resources is prestigious and worth adding to their CVs. We note that some respondents shared their frustration in the responses as they were offered, e.g., small company gifts for months of labor. As shown in Figure 6, 40% of the respondents, who spent 1 day to more than a month annotating data report negative experiences. That is, their work was not properly compensated, acknowledged, or recognised. This calls for a need to set guidelines and standards when using community or participatory research services.
@@ 5 Recommendations
While there has been a considerable amount of work on the ethics of best practices for building NLP and ML artefacts Bender and Friedman (2018); Leech et al. (2024); Mohammad (2022, 2023), our findings substantiate the fact that research on under-resourced languages presents additional challenges linked to the reliance on unconventional practices.
@@ 5.1 Center the People
Our findings show that various issues ought to be addressed early as research in the area lacks established standards and is subject to power differentials (e.g., some researchers reported their work being used by more notable individuals for fame). Many under-resourced languages are from what is called “the Global South” with a large number of them being spoken rather than written.
@@
Language is an important part of a population’s identity and technologies dealing with it have a direct impact on people’s lives. Past NLP work highlights how to engage with speakers and communities whose languages are in question Bird (2020); Bird and Yibarbuk (2024); Bird (2024); Cooper et al. (2024); Ramponi (2024). We further reinforce this argument with our findings, i.e., most researchers researchers report marked limitations in their languages (see Figure 4). Hence, when a researcher reaches out to a group of people with little background knowledge of their culture or language, one needs to approach these problems from the perspective of the community in question Bird (2022). The question of who is exactly served needs to be addressed early on to avoid any misconception of perceived needs for language technologies.
@@
In addition to the large percentage of our survey respondents who reported not being properly credited for their labor (Figure 5), there were cases of emotional manipulation (e.g., making emotional arguments such as how one’s labor will help the speakers of the language and that is compensation enough). One has to set rules and expectations with clear communication on the purpose of a given research project. For instance, when dealing with online participatory research communities for data collection and annotation, extra care needs to be shown and benevolent prejudice such as depicting oneself as a savior of a local community Bird (2022) must be avoided. Companies and research labs relying on communities for annotation and data creation need to properly compensate the contributors by, e.g., drawing legal contracts. The question of who is annotating what has to be addressed as well. The scarcity of annotators can lead to poor choices as very often, native speakers cannot be found online easily which has led to researchers choosing people from associated regions—people who do not necessarily speak the language variant in question. This results in a problematic overgeneralisation that puts different languages under the same umbrella simply because they have one or a small set of attributes in common. This often results in potentially over-simplistic solutions. For instance, variations of Arabic differ considerably but numerous research projects have treated entire regions, such as North Africa, as a monolith (e.g., to appear to have more data).
@@ 5.2 Be Fair: Give Credit where Credit is Due
Our findings show the unfortunate trend of data workers and NLP researchers suffering from a lack of recognition, especially those who are part of online communities (Figure 3). A needed follow-up work would be extensive fieldwork with the various online communities. Hence, our recommendation is a call for action on the setup of fair and comprehensive practices in participatory research projects, while taking power differentials into account. That is, existing authorship standardsneed to be followed and discussed prior to the start of a project as to whether a data worker should be listed as an author. This is particularly critical for junior researchers who substantially contribute to the resource construction and may assume that they will have a leading role in the paper. Moreover, proper financial compensation needs to be provided for annotators. Ideally, a resource paper should provide proof that the annotators were paid and treated fairly as recommended by Rogers et al. (2021).
@@ 5.3 Choose the Jargon Carefully and Be Aware of False Generalisations
As previously discussed in 5.1, it is important to embrace social awareness and avoid grouping people from colonial and Western perspectives Bird (2020, 2022); Held et al. (2023). Hence, one can avoid dismissive and outdated terms and classifications, e.g., “the rest of the world”. Note also that The World’s Values Survey classification Haerpfer and Kizilova (2012), which is often used in NLP papers (e.g., Santy et al. (2023)), presents an orientalist view of the world Said (1977). It has clear flaws such as including Christian-majority countries (e.g., Ethiopia, Rwanda) in a so-called "African-Islamic" category and groups of countries that have little in common in one category (e.g., Kyrgyzstan and Tunisia), leading to misrepresentations and enforcing stereotyping.
@@ 5.4 Set Fair and Realistic Expectations
As pointed out by Doğruöz and Sitaram (2022), tools for low-resource languages are often perceived as scaled-down versions of high-resource ones. Adding to previous work elaborating on what this may mean to the speakers Bird (2022); Markl et al. (2024), we focus on the impact of setting these expectations for researchers and practitioners. That is, they may be expected to build models similar to those built for high-resource languages, i.e., tackling the same NLP tasks, and performing extremely well. However, this can be unrealistic for various reasons such as the users’ needs Blaschke et al. (2024), the language’s specific features Bird and Yibarbuk (2024), and the lack of funding linked to the “prestige” of the language as reported by our respondents and similarly discussed by Mihalcea et al. (2024) in the context of LLM research.
@@
Joshi et al. (2020) conducted a survey on the state of NLP in various languages. They found that people do not necessarily want the tools that researchers think they need. However, we should not be limiting what NLP research on low- to mid-resource languages should be about. The most challenging part is to find a balance between what technologies local communities need and the research explored by the NLP community. This balance could be reinforced by continued communication with local communities Bird and Yibarbuk (2024); Lent et al. (2022); Mager et al. (2023); Cooper et al. (2024).
@@
Just because automatic systems obtain high scores on some English datasets (often making use of massive amounts of English data) does not mean that the problem is solved generally, or that working on that problem in other languages is no longer interesting or valuable. Working with each new language presents many new challenges, e.g., a rich morphology or the presence of tones Adebara and Abdul-Mageed (2022). Therefore, such work should not be undervalued by calling it “a replication” study.
@@ 5.5 Check the Source Even if the Language is Low-resource
Due to the limited amount of online data available for some languages, there is a tendency to use any available online source to build resources for these languages without examining the ethical implications or the appropriateness of the source. While it is typically easier to use religious texts, song lyrics, or movie subtitles, these should be carefully considered Hutchinson (2024); Mager et al. (2023). For instance, song lyrics are not representative of daily communication Mayer et al. (2008) since, e.g., they often rhyme and contain profanities. Additionally, using religious texts without a thorough inspection of potential implications can lead to misrepresentations (e.g., considering some cultures to be highly religious) Mager et al. (2023). Further, some work turns to synthetic data generated using machine translation and LLMs when these show clear limitations, especially in multicultural settings Hershcovich et al. (2022)(see Section 4.2.1). Research from other disciplines, even tightly related, such as linguistics Turner (2023), can help us choose adequate and suitable data sources.
@@ 6 Conclusion
We present insights from NLP researchers and practitioners working on under-served languages. We discuss limitations and research practices in the field, then provide actionable recommendations on how to address the reported issues while remaining fair to data workers. Our work is the first to document NLP researchers and workers’ experiences.
Limitations
@@ 7 Acknowledgments
Many thanks to Joanne Boisson, Michael Schlichtkrull, Sowmya Vajjala, and Evani Radiya-Dixit for their helpful feedback.
@@ A.2 Languages
The full list of the languages that our respondents have worked is included in the following. Note that participants could work on more than one language. Some may have also conducted research for high-resource languages and the respondents may or may not speak the language(s).